# On discount functions for economic model predictive control without terminal conditions

### v1

@@ 1 Introduction
Economic model predictive control (E-MPC) has gained popularity due to its ability to directly optimize a certain performance criterion while satisfying constraints for nonlinear systems [4], [5], [8], [10]. In E-MPC, the cost to be minimized by solving an optimal control problem could represent, e.g., the production capacity or energy demand in process control [4]. The optimal operating behavior in such a setting is not pre-specified as in tracking MPC, but implicitly specified by the interplay of the system, the stage cost, and the constraints. Thus, the optimal operation can result in a steady state (e.g., [1], [2], [6], [7]), a periodic orbit(e.g., [9][13][14][15]), or an even more general trajectory (e.g., [3][11]). To give guarantees on stability and performance, terminal conditions are used in [1] for steady state and in [14] for periodic optimal operation. In [14], however, the stability is only proven for one specific phase on the orbit. Moreover, implementing terminal conditions requires significant design effort including knowledge of the optimal behavior, which is why terminal conditions are often avoided in practice. For E-MPC without terminal conditions in scenarios where the optimal operating behavior is a steady state, [6] and [7] provide practical asymptotic stability under usual controllability and dissipativity assumptions. However, if the optimal operating behavior is a periodic orbit and the cost is not constant on the orbit [9], this MPC scheme could fail to approach optimal average performance even for an arbitrarily long prediction horizon as seen in examples in [15, Ex. 6] or [13, Ex. 4, Ex. 5]. To overcome this problem, the authors of [13] propose a $p^{\star}$-step E-MPC scheme, where $p^{\star}$ is the period length of the optimal orbit. However, this setup requires system- and cost-specific knowledge of the orbit length and can only provide asymptotic convergence but no stability guarantees. As a remedy, [15] introduced a linearly discounted stage cost function and provided guarantees on asymptotic average performance and practical asymptotic stability. This means that any arbitrarily small neighborhood of the optimal orbit can be asymptotically stabilized if the prediction horizon length is long enough. However, for some examples, extremely long prediction horizons may be necessary to achieve the desired performance [15, Ex. 19]. In this work, we investigate different discount functions for which we can get the same qualitative guarantees as in [15] and show in numerical examples that even for short prediction horizons the performance can be improved significantly. In particular, we prove stability for a large class of discount functions, thereby enabling the control engineer to use the discount function as a tuning variable. Further, it follows from this result that the asymptotic average performance is optimal up to an error that vanishes with growing prediction horizons. Our numerical examples show the performance of various discounts and we observe in these examples that the linear discount from [15] can be significantly outperformed if, for example, only the second half of the horizon is linearly discounted.
@@
Outline. In Section 2, we introduce the problem setup and the discounted E-MPC scheme. In Section 3, we present the main result on practical asymptotic stability, while the technical parts of the proof are moved to the Appendix A. Before drawing our conclusion in Section 5, we supplement this work in Section 4 with numerical examples to compare the performance of different discounts.
@@
Notation. In this work, $\mathbb{N}$ denotes the set of naturals and $\mathbb{R}$ the set of real numbers, whereas $\mathbb{R}_{0}^{+}$ denotes the set of all non-negative reals. Integers in the interval $[a,\,b]$ are denoted by $\mathbb{I}_{[a,b]}$ for $a\leq b$. The modulo operator $[k]_{p}$ returns the remainder when dividing $k\in\mathbb{N}$ by the natural $p\geq 1$. The floor-operator $\lfloor x\rfloor$ returns the largest integer that is smaller than or equal to $x\in\mathbb{R}$ and respectively $\lceil x\rceil$ as the ceiling-operator denotes the smallest integer that is larger than or equal to $x$. The function class $\mathcal{K}$ contains functions $\alpha:\mathbb{R}_{0}^{+}\to\mathbb{R}_{0}^{+}$, which are continuous and strictly increasing with $\alpha(0)=0$. It is $\alpha\in\mathcal{K}_{\infty}$ if $\alpha$ is additionally unbounded. We denote $\delta\in\mathcal{L}$ if $\delta:\mathbb{N}\to\mathbb{R}_{0}^{+}$ is non-increasing and $\lim_{k\to\infty}\delta(k)=0$. Further, $\gamma\in\mathcal{KL}$ denotes functions $\gamma:\mathbb{R}_{0}^{+}\times\mathbb{N}\to\mathbb{R}_{0}^{+}$ with the property $\gamma(\cdot,\,t)\in\mathcal{K}$ and $\gamma(r,\,\cdot)\in\mathcal{L}$.
@@ 2 Problem setup
We consider the same problem setup as in [15], that is a nonlinear discrete-time system
$x(k+1)=f(x(k),\,u(k)),\qquad x(0)=x_{0}$ (1)
with states and inputs subject to the constraints $x(k)\in\mathbb{X}\subseteq\mathbb{R}^{n}$ and $u(k)\in\mathbb{U}\subseteq\mathbb{R}^{m}$. The solution of (1) at time step $k$ resulting from a given input sequence $u\in\mathbb{U}^{K}$ of length $K\in\mathbb{N}$ and initial condition $x_{0}\in\mathbb{X}$ is denoted by $x_{u}(k,x_{0})$. By $\mathbb{U}^{K}(x_{0})$ we denote the set of feasible input sequences at $x_{0}\in\mathbb{X}$ of length $K$, i.e., $u\in\mathbb{U}^{K}$ with $x_{u}(k,x_{0})\in\mathbb{X}$ for all $k\in\mathbb{I}_{[0,K]}$. We aim to operate the system optimally concerning a stage cost $\ell:\mathbb{X}\times\mathbb{U}\to\mathbb{R}$ which is not required to be positive definite with respect to a desired setpoint and hence can be chosen according to an economic criterion. Define the accumulated cost for the finite horizon $T$ by
$J_{T}(x,\,u):=\sum_{k=0}^{T-1}\ell(x_{u}(k,\,x),\,u(k)),$ (2)
then the control objective is to minimize the asymptotic average performance
$J_{\infty}^{\mathrm{av}}(x,\,u):=\limsup_{T\to\infty}\frac{1}{T}J_{T}(x,\,u)$ (3)
and to stabilize the optimal operating behavior. We consider scenarios where the optimal operating behavior is a periodic orbit.
@@
A $p$-tuple $\Pi\in(\mathbb{X}\times\mathbb{U})^{p}$, $p\in\mathbb{N}$ is called a $p$-periodic orbit, if its projection $\Pi_{\mathbb{X}}$ onto $\mathbb{X}^{p}$ satisfies
$\Pi_{\mathbb{X}}([k+1]_{p})=f(\Pi(k))$ (4)
for all $k\in\mathbb{I}_{[0,p-1]}$. A $p$-periodic orbit $\Pi$ is called minimal, if $\Pi_{\mathbb{X}}(k)=\Pi_{\mathbb{X}}(j)\Rightarrow k=j$ for all $k,j\in\mathbb{I}_{[0,p-1]}$. The distance of a pair $(x,u)\in\mathbb{X}\times\mathbb{U}$ to the orbit $\Pi$ is defined by $\|(x,u)\|_{\Pi}:=\inf_{k\in\mathbb{I}_{[0,p-1]}}\|(x,u)-\Pi(k)\|$. The set of all feasible $p$-periodic orbits is denoted by $S_{\Pi}^{p}$. The average cost at $\Pi\in S_{\Pi}^{p}$ is defined by $\ell^{p}(\Pi):=\frac{1}{p}\sum_{k=0}^{p-1}\ell(\Pi(k))$. If a feasible $p^{\star}$-periodic orbit ${\Pi^{\star}}$ satisfies
$\ell^{p^{\star}}({\Pi^{\star}})=\inf_{p\in\mathbb{N},\,\Pi\in S_{\Pi}^{p}}\ell^{p}(\Pi)=:\ell^{\star},$ (5)
then ${\Pi^{\star}}$ is called an optimal periodic orbit and $p^{\star}$ is called an optimal period length.
To derive our results, we require the following assumptions.
@@
There exists an optimal periodic orbit $\Pi^{\star}$ of length $p^{\star}$ and it is minimal. There exists a continuous storage function $\lambda:\mathbb{X}\to\mathbb{R}$ and a function $\underline{\alpha}_{\tilde{\ell}}\in\mathcal{K}_{\infty}$, such that for all $x\in\mathbb{X}$ and $u\in\mathbb{U}^{1}(x)$ it holds that
~$\tilde{\ell}(x,u):=\ell(x,u)-\ell^{\star}+\lambda(x)-\lambda(f(x,u))\geq\underline{\alpha}_{\tilde{\ell}}(\lVert(x,u)\rVert_{\Pi^{\star}}).$ (6)
The function $\tilde{\ell}$ is called the rotated stage cost. The supply rate in the strict dissipativity of Assumption 1 is $\ell(x,u)-\ell^{\star}$.
Assumption 2 (Continuity, compactness, control invariance)
The functions $f$ and $\ell$ are continuous, $\ell$ is non-negative and the constraint set $\mathbb{X}\times\mathbb{U}$ is compact. Further, $\mathbb{X}$ is control invariant for system (1), i.e., for all $x\in\mathbb{X}$ there exists $u\in\mathbb{U}^{1}(x)$.
Due to Assumption 2, ¯$\bar{\ell}:=\sup_{(x,u)\in\mathbb{X}\times\mathbb{U}}\ell(x,u)$ and $\bar{\lambda}:=\sup_{x\in\mathbb{X}}|\lambda(x)|$ are finite.
@@
There exists $\kappa>0,\ M_{1}\in\mathbb{N},\ \alpha_{c}\in\mathcal{K}_{\infty}$ such that for any $z\in\Pi_{\mathbb{X}}^{\star}$ and all $x,\,y\in\mathbb{X}$ with $\lVert x-z\rVert\leq\kappa$ and $\lVert y-z\rVert\leq\kappa$ a control input sequence $u\in\mathbb{U}^{M_{1}}(x)$ exists satisfying $x_{u}(M_{1},\,x)=y$ and
$\lVert(x_{u}(k,\,x),\,u(k))\rVert_{\Pi^{\star}}\leq\alpha_{c}(\mathrm{max}\{\lVert x\rVert_{\Pi_{\mathbb{X}}^{\star}},\,\lVert y\rVert_{\Pi_{\mathbb{X}}^{\star}}\}),\quad\forall k\in\mathbb{I}_{[0,\,M_{1}-1]}.$ (7)
@@
These assumptions are standard in the context of E-MPC (compare [7][9][13]). In particular, they are the same as the ones required to guarantee practical asymptotic stability of the linearly discounted economic MPC scheme from [15] with one exception: Control invariance of $\mathbb{X}$ is additionally required to simplify the exposition of our results. Analogous to [15, Section 4], the results could be shown as well for all points from which $\Pi^{\star}$ can be reached in a limited time even if control invariance of $\mathbb{X}$ is not given.
@@ 2.1 Discounted Economic MPC
$V_{N}^{\beta}(x)=\min_{u\in\mathbb{U}^{N}(x)}J_{N}^{\beta}(x,u)=\min_{u\in\mathbb{U}^{N}(x)}\sum_{k=0}^{N-1}{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\ell\left(x_{u}(k,x),u(k)\right).$ (8)
Assumption 2 guarantees that for all $x\in\mathbb{X}$ and all $N\in\mathbb{N}$ a (possibly non-unique) optimal control sequence $u_{x}^{\star}$ exists, i.e, $V_{N}^{\beta}(x)=J_{N}^{\beta}(x,u_{x}^{\star})$. For convenience of notation, we will also use $x_{x}^{\star}(k):=x_{u_{x}^{\star}}(k,x)$. Although this lucid notation hides it, both $u_{x}^{\star}$ and $x_{x}^{\star}$ depend on the prediction horizon length $N$ and the chosen discount function $\beta$. In the same way, we define $\tilde{J}_{N}^{\beta}$, $\tilde{V}_{N}^{\beta}$, $\tilde{u}_{x}^{\star}$, and $\tilde{x}_{x}^{\star}$ for the case when we replace $\ell$ by $\tilde{\ell}$ from (6). The MPC feedback we apply is $\mu_{N}^{\beta}(x)=u_{x}^{\star}(0)$. The discount function $\beta:[0,1]\to[0,1]$ defines a uniform shape of the discount for all horizon lengths $N$. This is crucial as the following example shows.
@@
Let us consider the example depicted in Fig. 1, which is from [13, Example 4] and has been used in [15, Example 6] as well. As shown therein, a non-discounted E-MPC scheme fails to approach the optimal orbit for every odd prediction horizon, whereas a linearly discounted scheme with $\beta^{\text{lin}}(\frac{k}{N})=1-\frac{k}{N}$ succeeds for all prediction horizon lengths $N\geq 2$. Here, we want to consider a slight modification of this linear discount $\beta_{N}(k)=1-\frac{2\lceil k/2\rceil}{N}$, which has $\beta_{N}(k-1)=\beta_{N}(k)$ for all $k\in 2\mathbb{N}$. We use the modified notation $\beta_{N}(k)$ as this discount is not uniform, i.e., there exists no function $\beta:[0,1]\to[0,1]$ such that $\beta_{N}(k)={\textstyle\beta\big{(}\frac{k}{N}\big{)}}$ for all $N\in\mathbb{N}$ and $k\in\mathbb{I}_{[0,N-1]}$. The question in this example is whether the optimal solution for horizon $N$ goes directly to the optimal orbit or waits one step before going there. When comparing these strategies, we need to compare only the cost for $k\in\{0,1,2\}$ as for $k\geq 3$ both strategies are on the optimal orbit and contribute the same cost as $\beta_{N}(k-1)=\beta_{N}(k)$ for all $k\in 2\mathbb{N}$ and the horizon length is effectively always odd as for even $N$ we have $\beta_{N}(N-1)=0$. Waiting has a cost of $1+\frac{N-2}{N}$ in these first $3$ steps and is hence the optimal strategy for all $N\in\mathbb{N}$ as going directly to the orbit costs $1+1.5\frac{N-2}{N}$. MPC applies only the first step of the optimal strategy and ends up waiting forever.
When using a uniform discount, this problem cannot occur.
11110011-1- $\ell(x,u)=1$$u=0$$\ell(x,u)=0$$u=1$$u=0$$\ell(x,u)=1.5$
$\ell(x,u)=1$ xxx$+\varepsilon$
$u=-1$
Figure 1: Illustration of the states $x$ (nodes) and feasible transitions (edges) with corresponding input $u$ and cost $\ell$ in Example 2.3. This diagram is taken from [15].
To show that the discounted E-MPC scheme achieves practical asymptotic stability of the optimal orbit $\Pi^{\star}$ we need the following assumptions on $\beta$.
@@
The discount function $\beta:[0,1]\to[0,1]$ is piecewise continuously differentiable, Lipschitz-continuous with Lipschitz constant $L_{\beta}$, non-increasing, and satisfies $\beta(0)=1$, $\beta(1)=0$, and $\mathrm{lim\,inf}_{\xi\to 1}\beta^{\prime}(\xi)>0$.
@@
Discount functions that satisfy Assumption 5 are, for example, the linear discount $\beta^{\text{lin}}(\xi)=1-\xi$ from [15], the half-non-discounted-half-linear discount $\beta^{\text{half-lin}}(\xi)=\min\{1,2-2\xi\}$, the polynomial discount $\beta^{\mathrm{poly}}(\xi)=1-\xi^{q}$ with degree $q\in\mathbb{N}$, and many more. The non-discounted discount $\beta^{\text{non}}(\xi)=1$ for all $\xi\in[0,1]$ does not satisfy $\beta(1)=0$.

### v2

# Lukas Schwenkel, Daniel Briem, Matthias Müller, Frank Allgöwer

@@ 1 Introduction
Economic model predictive control (E-MPC) has gained popularity due to its ability to directly optimize a certain performance criterion while satisfying constraints for nonlinear systems [4], [5], [9], [11]. In E-MPC, the cost to be minimized by solving an optimal control problem could represent, e.g., the production capacity or energy demand in process control [4]. The optimal operating behavior in such a setting is not pre-specified as in tracking MPC, but implicitly specified by the interplay of the system, the stage cost, and the constraints. Thus, the optimal operation can result in a steady state (e.g., [1], [2], [7], [8]), a periodic orbit(e.g., [10][14][16][17]), or an even more general trajectory (e.g., [3][12]). To give guarantees on stability and performance, terminal conditions are used in [1] for steady state and in [16] for periodic optimal operation. In [16], however, the stability is only proven for one specific phase on the orbit. Moreover, implementing terminal conditions requires significant design effort including knowledge of the optimal behavior, which is why terminal conditions are often avoided in practice. For E-MPC without terminal conditions in scenarios where the optimal operating behavior is a steady state, [7] and [8] provide practical asymptotic stability under usual controllability and dissipativity assumptions. However, if the optimal operating behavior is a periodic orbit and the cost is not constant on the orbit [10], this MPC scheme could fail to approach optimal average performance even for an arbitrarily long prediction horizon as seen in examples in [17, Ex. 6] or [14, Ex. 4, Ex. 5]. To overcome this problem, the authors of [14] propose a $p^{\star}$-step E-MPC scheme, where $p^{\star}$ is the period length of the optimal orbit. However, this setup requires system- and cost-specific knowledge of the orbit length and can only provide asymptotic convergence but no stability guarantees. As a remedy, [17] introduced a linearly discounted stage cost function and provided guarantees on asymptotic average performance and practical asymptotic stability. This means that any arbitrarily small neighborhood of the optimal orbit can be asymptotically stabilized if the prediction horizon length is long enough. However, for some examples, extremely long prediction horizons may be necessary to achieve the desired performance [17, Ex. 19].
In this work, we investigate for which discount functions we can get the same qualitative guarantees as in [17] and show in numerical examples that even for short prediction horizons the performance can be improved significantly. In particular, we prove stability for a large class of discount functions, thereby enabling the control engineer to use the discount function as a tuning variable. Further, it follows from this result that the asymptotic average performance is optimal up to an error that vanishes with growing prediction horizons. Our numerical examples show the performance of various discounts and we observe in these examples that the linear discount from [17] can be significantly outperformed if, for example, only the second half of the horizon is linearly discounted. We emphasize that this work aims to solve the undiscounted infinite-horizon optimal control problem. This differs from related works on exponentially discounted economic MPC (e.g., [6][15], and references therein), which focus on solving the exponentially discounted infinite-horizon optimal control problem.
@@
Outline. Section 2 introduces the problem setup and the discounted E-MPC scheme. Section 3 contains the main results, while the technical parts of the proof are moved to the Appendix A. The numerical examples are presented in Section 4 and our conclusion are drawn in Section 5.
@@
Notation. In this work, $\mathbb{N}$, $\mathbb{R}$, and $\mathbb{R}_{0}^{+}$ denote the sets of naturals, reals, and non-negative reals. Integers in the interval $[a,\,b]$ are denoted by $\mathbb{I}_{[a,b]}$ for $a\leq b$. The modulo operator $[k]_{p}$ returns the remainder when dividing $k\in\mathbb{N}$ by the natural $p\geq 1$. The floor-operator $\lfloor x\rfloor$ returns the largest integer that is smaller than or equal to $x\in\mathbb{R}$ and respectively $\lceil x\rceil$ as the ceiling-operator denotes the smallest integer that is larger than or equal to $x$. The function class $\mathcal{K}$ contains functions $\alpha:\mathbb{R}_{0}^{+}\to\mathbb{R}_{0}^{+}$, which are continuous and strictly increasing with $\alpha(0)=0$. It is $\alpha\in\mathcal{K}_{\infty}$ if $\alpha$ is additionally unbounded. We denote $\delta\in\mathcal{L}$ if $\delta:\mathbb{N}\to\mathbb{R}_{0}^{+}$ is non-increasing and $\lim_{k\to\infty}\delta(k)=0$. Further, $\gamma\in\mathcal{KL}$ denotes functions $\gamma:\mathbb{R}_{0}^{+}\times\mathbb{N}\to\mathbb{R}_{0}^{+}$ with the property $\gamma(\cdot,\,t)\in\mathcal{K}$ and $\gamma(r,\,\cdot)\in\mathcal{L}$.
@@ 2 Problem setup
We consider a nonlinear discrete-time system
$x(k+1)=f\big{(}x(k),u(k)\big{)},\qquad x(0)=x_{0}$ (1)
with state and input constraints $x(k)\in\mathbb{X}\subseteq\mathbb{R}^{n}$ and $u(k)\in\mathbb{U}\subseteq\mathbb{R}^{m}$. The solution of (1) at time step $k$ resulting from an input sequence $u\in\mathbb{U}^{K}$ of length $K\in\mathbb{N}$ and initial condition $x_{0}\in\mathbb{X}$ is denoted by $x_{u}(k,x_{0})$. The input sequence $u\in\mathbb{U}^{K}$ is feasibly, denoted by $u\in\mathbb{U}^{K}(x_{0})$, if $x_{u}(k,x_{0})\in\mathbb{X}$ for all $k\in\mathbb{I}_{[0,K]}$. We aim to operate the system optimally concerning a (not necessarily positive definite) stage cost $\ell:\mathbb{X}\times\mathbb{U}\to\mathbb{R}$, in particular, we want to minimize the asymptotic average performance
$J_{\infty}^{\mathrm{av}}(x,u):=\limsup_{T\to\infty}\frac{1}{T}J_{T}(x,u)\ \ \text{ with }\ \ J_{T}(x,u):=\sum_{k=0}^{T-1}\ell\big{(}x_{u}(k,x),u(k)\big{)}$ (2)
and to stabilize the optimal operating behavior. We assume that the optimal operating behavior is a periodic orbit, which is reachable and controllable.
@@
Let $\Pi_{\mathbb{X}}:\mathbb{I}_{[0,p-1]}\to\mathbb{X}$ and $\Pi_{\mathbb{U}}:\mathbb{I}_{[0,p-1]}\to\mathbb{U}$ for $p\in\mathbb{N}$. Then $\Pi=(\Pi_{\mathbb{X}},\Pi_{\mathbb{U}})$ is called a $p$-periodic orbit of system (1), if
$\Pi_{\mathbb{X}}\big{(}[k+1]_{p}\big{)}=f\big{(}\Pi_{\mathbb{X}}(k),\Pi_{\mathbb{U}}(k)\big{)}=f\big{(}\Pi(k)\big{)}$ (3)
for all $k\in\mathbb{I}_{[0,p-1]}$. A $p$-periodic orbit $\Pi$ is called minimal, if $\Pi_{\mathbb{X}}$ is injective. The distance of a point $(x,u)\in\mathbb{X}\times\mathbb{U}$ to the image of the function $\Pi$ is defined by $\|(x,u)\|_{\Pi}:=\min_{k\in\mathbb{I}_{[0,p-1]}}\|(x,u)-\Pi(k)\|$.
Assumption 1 (Optimal periodic orbit)
There exists a minimal $p^{\star}$-periodic orbit ${\Pi^{\star}}$ that satisfies for any $p\in\mathbb{N}$ and any $p$-periodic orbit $\Pi$ that
$\ell^{\star}:=\frac{1}{p^{\star}}\sum_{k=0}^{{p^{\star}}-1}\ell(\Pi^{\star}(k))\leq\frac{1}{p}\sum_{k=0}^{p-1}\ell(\Pi(k)).$ (4)
The following assumptions characterize the problem class for which we can provide performance and stability guarantees. These assumptions are standard in the context of periodic E-MPC (cf. [10], [14][16][17]).
@@
There exist $\underline{\alpha}_{\tilde{\ell}}\in\mathcal{K}_{\infty}$ and a continuous storage function $\lambda:\mathbb{X}\to\mathbb{R}$ such that for all $x\in\mathbb{X}$ and $u\in\mathbb{U}^{1}(x)$ we have
~$\tilde{\ell}(x,u):=\ell(x,u)-\ell^{\star}+\lambda(x)-\lambda\big{(}f(x,u)\big{)}\geq\underline{\alpha}_{\tilde{\ell}}(\lVert(x,u)\rVert_{\Pi^{\star}}).$ (5)
We call $\ell(x,u)-\ell^{\star}$ the supply rate and $\tilde{\ell}$ the rotated stage cost.
Assumption 3 (Continuity and compactness)
The functions $f$ and $\ell$ are continuous, and the constraint set $\mathbb{X}\times\mathbb{U}$ is compact. Hence, $\bar{\lambda}:=\sup_{x\in\mathbb{X}}|\lambda(x)|$ and ¯$\bar{\ell}:=\sup_{(x,u)\in\mathbb{X}\times\mathbb{U}}|\ell(x,u)|$ are finite.
@@
There exists $\kappa>0,\ M_{1}\in\mathbb{N},\ \alpha_{c}\in\mathcal{K}_{\infty}$ such that for all $z\in\Pi_{\mathbb{X}}^{\star}$ and $x,\,y\in\mathbb{X}$ with $\lVert x-z\rVert\leq\kappa$ and $\lVert y-z\rVert\leq\kappa$ a control input sequence $u\in\mathbb{U}^{M_{1}}(x)$ exists with $x_{u}(M_{1},x)=y$ and
$\big{\lVert}\big{(}x_{u}(k,x),u(k)\big{)}\big{\rVert}_{\Pi^{\star}}\leq\alpha_{c}\big{(}\mathrm{max}\big{\{}\lVert x\rVert_{\Pi_{\mathbb{X}}^{\star}},\lVert y\rVert_{\Pi_{\mathbb{X}}^{\star}}\big{\}}\big{)},\quad\forall k\in\mathbb{I}_{[0,\,M_{1}-1]}.$ (6)
@@
We assume finite-time reachability of ${\Pi^{\star}}$ for all $x\in\mathbb{X}$ for simplicity of our proofs, similar to [14]. As detailed in [17, Section 4] for the linearly discounted E-MPC scheme, this assumption can be relaxed to hold only on a subset of $\mathbb{X}$. A similar relaxation is also possible in the setting with general discounts, which is, however, beyond the scope of the present paper.
@@ 2.1 Discounted Economic MPC
$V_{N}^{\beta}(x)=\min_{u\in\mathbb{U}^{N}(x)}J_{N}^{\beta}(x,u)=\min_{u\in\mathbb{U}^{N}(x)}\sum_{k=0}^{N-1}{\textstyle\beta\left(\frac{k}{N}\right)}\ell\big{(}x_{u}(k,x),u(k)\big{)}.$ (7)
Assumption 3 guarantees that for all $x\in\mathbb{X}$ and all $N\in\mathbb{N}$ a (possibly non-unique) optimal control sequence $u_{x}^{\star}$ exists, i.e, $V_{N}^{\beta}(x)=J_{N}^{\beta}(x,u_{x}^{\star})$. For convenience of notation, we will also use $x_{x}^{\star}(k):=x_{u_{x}^{\star}}(k,x)$. Although this lucid notation hides it, both $u_{x}^{\star}$ and $x_{x}^{\star}$ depend on the prediction horizon length $N$ and the chosen discount function $\beta$. The MPC feedback we apply is $\mu_{N}^{\beta}(x)=u_{x}^{\star}(0)$. The discount function $\beta:[0,1]\to[0,1]$ defines a uniform shape of the discount for all horizon lengths $N$. This is crucial as the following example shows.
@@
Let us consider the example $x(t+1)=u(t)$ with only three feasible states $\mathbb{X}=\{-1,0,1\}$ and four feasible transitions $\{(-1,-1),(-1,0),(0,1),(1,0)\}\subseteq\mathbb{X}\times\mathbb{U}$. Thus, the only possible decision to make is at $x=-1$ whether to go to $x=0$ and alternate at the $2$-periodic orbit between $x=0$ and $x=1$ or whether to stay at $x=-1$. The cost values are $\ell(-1,-1)=1$, $\ell(-1,0)=1$, $\ell(0,1)=0$, $\ell(1,0)=1.5$ and hence the system is optimally operated at the periodic orbit with average cost $\ell^{\star}=0.75$, which is better than staying at $x=-1$ for cost $1$. However, when considering optimal trajectories of odd length $N$, it is better to wait one step at $x=-1$ for cost $1$ and then approach the orbit rather than approaching it directly and ending with the step of cost $1.5$ on the orbit. It was discovered in [14, Example 4] that this phenomenon causes an undiscounted E-MPC scheme with odd prediction horizon length to wait at $x=-1$ forever. In [17, Example 6] it was shown that linear discounts $\beta^{\text{lin}}(\frac{k}{N})=1-\frac{k}{N}$ solve this problem and that a linearly discounted E-MPC goes directly to the optimal orbit for all prediction horizons $N\geq 2$. Let us consider a slight modification of this linear discount $\beta_{N}(k-1)=\beta_{N}(k)=\beta^{\text{lin}}(\frac{k}{N})$ for all $k\in 2\mathbb{N}$. We introduced the notation $\beta_{N}(k)$ as this discount is not uniform, i.e., there exists no function $\beta:[0,1]\to[0,1]$ such that $\beta_{N}(k)={\textstyle\beta\left(\frac{k}{N}\right)}$ for all $N\in\mathbb{N}$ and $k\in\mathbb{I}_{[0,N-1]}$. Due to the special shape of the discount, when comparing the strategy of waiting one step with the one of directly going to the orbit, we need to compare only the cost for $k\in\{0,1,2\}$ as for $k\geq 3$ both strategies are on the optimal orbit and contribute the same cost as $\beta_{N}(k-1)=\beta_{N}(k)$ for all $k\in 2\mathbb{N}$ and the horizon length is effectively always odd as for even $N$ we have $\beta_{N}(N-1)=0$. Waiting one step has a cost of $1+\frac{N-2}{N}$ in these first $3$ steps and is hence the optimal strategy for all $N\in\mathbb{N}$ as going directly to the orbit costs $1+1.5\frac{N-2}{N}$. As MPC applies only the first step of the optimal strategy it ends up waiting forever. When using a uniform discount, this problem cannot occur.
The following assumption confines the choice of the discount function $\beta$. This is required to guarantee practical asymptotic stability of the discounted E-MPC.
@@
The discount function $\beta:[0,1]\to[0,1]$ is non-increasing, piecewise continuously differentiable, Lipschitz-continuous with Lipschitz constant $L_{\beta}$, and satisfies $\beta(0)=1$, $\beta(1)=0$, and $\mathrm{lim\,sup}_{\xi\to 1}\beta^{\prime}(\xi)<0$.
The technical assumption $\mathrm{lim\,sup}_{\xi\to 1}\beta^{\prime}(\xi)<0$ is made to simplify our proofs and does not pose a relevant constraint in practice.
@@
Discount functions that satisfy Assumption 6 are, for example, the linear discount $\beta^{\text{lin}}(\xi)=1-\xi$ from [17], the half-undiscounted-half-linear discount $\beta^{\text{half-lin}}(\xi)=\min\{1,2-2\xi\}$, the polynomial discount $\beta^{\mathrm{poly}}(\xi)=1-\xi^{q}$ with degree $q\in\mathbb{N}$, and many more. The undiscounted discount $\beta^{\text{un}}(\xi)=1$ for all $\xi\in[0,1]$ does not satisfy $\beta(1)=0$.